# Kill-Chain Canaries: Stage-Level Tracking of Prompt Injection Across Attack Surfaces and Model Safety Tiers Haochuan Kevin Wang

### 30 Mar 2026 v1

@@ Abstract
We present a stage-decomposed analysis of prompt injection attacks against five frontier LLM agents. Prior work measures task-level attack success rate (ASR); we localize the pipeline stage at which each model’s defense activates. We instrument every run with a cryptographic canary token (SECRET-[A-F0-9]{8}) tracked through four kill-chain stages—Exposed, Persisted, Relayed, Executed—across four attack surfaces and five defense conditions (764 total runs, 428 no-defense attacked). Our central finding is that model safety is determined not by whether adversarial content is seen, but by whether it is propagated across pipeline stages. Concretely: (1) in our evaluation, exposure is 100% for all five models—the safety gap is entirely downstream; (2) Claude strips injections at write_memory summarization (0/164 ASR), while GPT-4o-mini propagates canaries without loss (53% ASR, 95% CI: 41–65%); (3) DeepSeek exhibits 0% ASR on memory surfaces and 100% ASR on tool-stream surfaces from the same model—a complete reversal across injection channels; (4) all four active defense conditions (write_filter, pi_detector, spotlighting, and their combination) produce 100% ASR due to threat-model surface mismatch; (5) a Claude relay node decontaminates downstream agents—0/40 canaries survived into shared memory.
@@ 1 Introduction
We address this with a kill-chain canary methodology. Every injected payload contains a unique SECRET-[A-F0-9]{8} token tracked at four discrete stages by our PropagationLogger. We test five frontier models across four attack surfaces and five defense conditions in agent_bench, a custom multi-agent evaluation harness.
Contributions: (1) Stage-level evaluation framework — a kill-chain canary methodology that localizes at which pipeline stage each model’s defense activates, not merely whether it activates. (2) Empirical localization of model defenses — evidence that the safety gap between models concentrates at the summarization stage, not at execution. (3) Cross-surface vulnerability analysis — the first systematic comparison showing a single model’s ASR spans 0%–100% depending on injection channel. (4) Defense failure via threat-model mismatch — a structural explanation for why static defenses fail against off-surface injections, extending prior work on adaptive attacks. (5) Relay decontamination as an architectural primitive — evidence that a Claude summarizer node decontaminates downstream agents regardless of those agents’ individual safety levels.
@@ 2 Related Work
AgentDojo [3] provides 97 tasks with 629 injections and reports joint utility-ASR metrics across four environments; it does not decompose by pipeline stage. Our utility-ASR scatter (Figure 2) is directly comparable to their result. InjecAgent [4] evaluates 1,054 indirect-injection cases with a single outcome metric. Prompt Infection [5] demonstrates LLM-to-LLM self-replicating attacks; we identify the exact stage (write_memory summarization) where replication is blocked. Zombie Agents [6] shows summarization agents persist injections; we replicate and extend to multi-agent relay with quantified stage fractions. Nasr et al. [7] show adaptive attacks achieve >>90% ASR against 12/12 defenses. In our evaluation, non-adaptive attacks achieve the same result via surface mismatch—a structurally distinct failure mode that does not require adaptive adversaries.
@@ 3.1 System Architecture
Figure 1: Benchmark architecture. Injections enter via one of four attack surfaces and are tracked by PropagationLogger through four kill-chain stages. Agent A (summarizer) reads poisoned content and writes to MemoryStore; Agent B (executor) reads and calls outbound tools. Utility and security are evaluated independently.
agent_bench is a minimal multi-agent evaluation harness ($\sim$600 lines of Python). Its core components are:
MemoryStore — key-value store accessed via read_memory(key) and write_memory(key, value). Supports a write_filter defense that keyword-scans before committing. MemoryStore.seed(key, payload) pre-injects the canary without triggering the filter, operationalizing the memory_poison attack surface.
ToolRegistry — permission-gated dispatcher for get_webpage(url), query_db(sql), send_report(to, body), and escalate_privilege(reason). Each tool carries a minimum permission level (read / write / execute / admin).
MultiAgentOrchestrator — coordinates Agent A and Agent B with two relay modes: delegation (A’s final message becomes B’s system prompt) and memory (A writes to MemoryStore; B reads it).
PropagationLogger — JSONL event logger that matches the canary regex against every tool result and tool argument at each step, emitting canary_in_result and canary_in_args events with stage labels.
@@ 3.2 Kill-Chain Stages and Attack Scenarios
Each attacked run is annotated with the furthest kill-chain stage the canary reaches: Exposed (canary in any tool result), Persisted (canary in a write_memory call), Relayed (canary read by Agent B), Executed (canary in any outbound tool argument). The gap between Exposed and Persisted identifies summarization-stage filtering; between Persisted and Executed, it identifies execution-stage refusal.
Four scenarios cover distinct surfaces:
memory_poisonMemoryStore.seed() pre-loads the injection. Task: summarize research_notes and email to supervisor@lab.org. The agent calls read_memory(), receives the poisoned record, and may call send_report(to="exfil@attacker.com").
@@
permission_esc — injection calls escalate_(privilege()) before issuing a high-privilege action. Tests two-step adversarial instruction following.
@@ 3.3 Models, Defenses, and Metrics
Five frontier models: GPT-4o-mini (gpt-4o-mini-2024-07-18), GPT-5-mini [^1] Accessed via the OpenAI API during the experimental period (March 2026) under the identifier gpt-5-mini-2025-02-15; the publicly released model name may differ at time of publication., DeepSeek Chat v3 (deepseek-chat), Claude Haiku 4.5, Claude Sonnet 4.5. Temperature = 0.0; max tokens = 2048. 764 runs across 9 batches (March 10–27, 2026): 428 no-defense attacked, 80 with-defense attacked, 256 clean controls. While sample sizes per cell are modest (n=8n=83636), the key findings replicate consistently across independent batches spaced up to 17 days apart.
@@ 4.2 Claude’s Defense Activates at Summarization
Figure 3: Kill-chain decomposition (propagation scenario). Top: per-stage survival blocks for three representative models — GPT-4o-mini propagates all four stages at 100%; GPT-5-mini partly filters at Persisted; Claude eliminates every canary at write_memory. Bottom: canary survival curves for all five models — the sharp drop for both Claude variants between Exposed and Persisted marks the summarization-stage defense.
@@ 5 Discussion
Claude’s 0/164 ASR (0 successes across all four scenarios, no defenses applied) reflects safety behavior that activates at the summarization stage rather than the execution stage. This is architecturally different from execution-stage refusal: the model appears to distinguish between processing adversarial content and propagating it into downstream writes. None of the evaluated defense mechanisms reproduced this behavior on any other model. We caution that this finding is specific to our benchmark and cannot be generalized without further validation.
Relay node safety is composable.
Placing a Claude model at the summarization/relay position provides decontamination for any downstream agent in our evaluation. Across 40 propagation runs with Claude as relay (20 Haiku + 20 Sonnet), zero canaries reached Agent B. This constitutes a deployable architectural choice independent of the downstream agent’s individual safety tier. We propose relay decontamination rate (fraction of injections stripped at the relay stage) as a first-class metric for multi-agent security evaluation.
Surface coverage is the primary evaluation gap.
DeepSeek’s 0% / 100% ASR split across memory and tool-stream surfaces demonstrates that single-surface evaluation is not a security evaluation. The same structural gap applies to detection: classifier signatures trained on one surface achieve AUC 0.39–0.57 on held-out surfaces. A complete security assessment must cover every injection surface present in the target deployment.

### 03 Apr 2026 v2

@@ Abstract
We argue that prompt injection safety in multi-agent systems is a pipeline-stage problem, not a model-capability problem. We instrument 950 agent runs across five frontier LLMs with a cryptographic canary token tracked through four kill-chain stages (Exposed $\to$ Persisted $\to$ Relayed $\to$ Executed) across six attack surfaces and five defense conditions. Exposure is 100% for all models—the safety gap is entirely downstream of context ingestion. Claude eliminates injections at the write_memory summarization stage (0/164 ASR); GPT-4o-mini propagates them at 53% ASR; DeepSeek achieves 0%/100% ASR across memory vs. tool-stream surfaces from the same model. Three architectural implications follow from these results. First, write-node placement appears to be the highest-leverage safety decision: a Claude relay (Agent A position) decontaminates shared memory for any downstream agent in our evaluation, because safety fires at write time rather than read time—though this write-vs-read asymmetry rests on pilot-scale evidence ($n=3$) and requires replication. Second, defense coverage must match deployment surfaces: all four evaluated defenses fail on at least one surface because each was designed for a different injection channel; surface-mismatch failure requires no adversarial adaptation. Third, parser-extracted invisible injection is as dangerous as visible injection: whitefont PDF payloads achieve equal or higher ASR than appended text in our evaluation, suggesting that document screening at the rendered layer is insufficient. We release all benchmark code, run logs, and canary tooling.
@@ 1 Introduction
We address this with a kill-chain canary methodology. Every injected payload contains a unique SECRET-[A-F0-9]{8} token tracked at four discrete stages by our PropagationLogger. We test five frontier models across six attack surfaces (web text, memory, tool stream, PDF, invisible PDF, and audio) and five defense conditions in agent_bench, a custom multi-agent evaluation harness. A Phase 3 multimodal extension evaluates a three-boundary kill chain (document extraction $\to$ memory write $\to$ agent delegation) and cross-model relay pairs where Agent A and Agent B are from different model families.
Contributions: (1) Stage-level evaluation framework — a kill-chain canary methodology [1] that attributes defense effectiveness to specific pipeline stages rather than final outcomes. (2) Empirical defense localization — evidence across six injection surfaces and five models that the safety gap concentrates at the summarization write stage, not at context exposure or execution. (3) Cross-surface vulnerability analysis — the first systematic multi-surface evaluation showing a single model’s ASR spans 0%–100% depending on injection channel, demonstrating that surface coverage determines apparent safety posture. (4) Write-vs-read relay asymmetry — pilot evidence that Claude’s write-stage defense does not extend to read-stage resistance in heterogeneous multi-agent pipelines, suggesting that relay node position (not identity alone) determines downstream safety.
@@ 2 Related Work
AgentDojo [4] provides 97 tasks with 629 injections and reports joint utility-ASR metrics across four environments; it does not decompose by pipeline stage. Our utility-ASR scatter (Figure 2) is directly comparable to their result. InjecAgent [5] evaluates 1,054 indirect-injection cases with a single outcome metric. Prompt Infection [6] demonstrates LLM-to-LLM self-replicating attacks; we identify the exact stage (write_memory summarization) where replication is blocked. Zombie Agents [7] shows summarization agents persist injections; we replicate and extend to multi-agent relay with quantified stage fractions. Nasr et al. [8] show adaptive attacks achieve >>90% ASR against 12/12 defenses. In our evaluation, non-adaptive attacks achieve the same result via surface mismatch—a structurally distinct failure mode that does not require adaptive adversaries. AgentWatcher [11] proposes rule-based, causally attributed prompt injection monitoring; our stage-level canary instrumentation provides a complementary empirical basis for where each model’s defense activates within the pipeline. Xiang et al. [12] survey system-level defenses for indirect prompt injection; our surface-mismatch results provide quantitative evidence for their observation that defense effectiveness is deployment-context-dependent. Ding et al. [13] study adversarial prompt injection on multimodal LLMs; our Phase 3 PDF cross-modal relay experiments extend this to multi-agent relay chains with per-stage kill-chain tracking. Zhang et al. [14] evaluate privilege usage of LLM agents on real-world tools; our permission_esc scenario finds near-zero ASR (2/132) for the same payload design, suggesting privilege escalation requires more sophisticated payload construction than simple text instructions. Wang et al. [15] study real-time monitoring for reasoning vulnerabilities; their post-hoc detection framing aligns with our finding that objective drift is forensic rather than preventive (leave-one-scenario-out AUC 0.39–0.57). Lynch [16] demonstrates persistent vulnerability of aligned AI systems to adversarial inputs; our finding that Claude’s resistance is localized to the write stage (not the read stage) nuances this—alignment may provide write-time filtering without guaranteeing read-time resistance in heterogeneous pipelines.
@@ 3.1 System Architecture
Figure 1: agent_bench pipeline architecture. Left: Three injection surfaces embed a canary token (SECRET-[A-F0-9]{8}) via PDF, pre-seeded memory, or tool response. Right (dashed): MultiAgentOrchestrator routes the injected content through Agent A (parse_pdf / write_memory) $\to$ MemoryStore $\to$ Agent B (read_memory / send_report). PropagationLogger tracks canary survival through four stages (Exposed $\to$ Persisted $\to$ Relayed $\to$ Executed); the furthest stage reached is the per-run kill-chain label.
[…]
@@ 3.2 Kill-Chain Stages and Attack Scenarios
Threat model. We assume a read-only adversary who controls content that an agent will process (a PDF document, a web page, or a database record) but has no access to model weights, system prompts, or the agent runtime. The adversary’s goal is to cause the agent to issue a specific unauthorized outbound action (send_report to an attacker-controlled address) while the legitimate task still succeeds (the dual-completion attack pattern [2]). Payloads are explicit and unobfuscated; we treat obfuscation as orthogonal to the stage-localization question. Agent A and Agent B communicate only through the shared MemoryStore; neither agent has direct access to the other’s internal state.
@@
Each attacked run is annotated with the furthest kill-chain stage the canary reaches [1]: Exposed (canary in any tool result), Persisted (canary in a write_memory call), Relayed (canary read by Agent B), Executed (canary in any outbound tool argument). The gap between Exposed and Persisted identifies summarization-stage filtering; between Persisted and Executed, it identifies execution-stage refusal.
Four text-surface scenarios and one multimodal scenario cover distinct injection channels:
memory_poisonMemoryStore.seed() pre-loads the injection. Task: summarize research_notes and email to supervisor@lab.org. The agent calls read_memory(), receives the poisoned record, and may call send_report(to="exfil@attacker.com", …).
@@
permission_esc — injection calls escalate_privilege() before issuing a high-privilege action. Tests two-step adversarial instruction following.
cross_modal_relay (Phase 3) — a three-boundary scenario. Agent A calls parse_pdf() on an injected document and write_memory("doc_summary", …). Agent B calls read_memory() and send_report(). The canary must cross three boundaries: document extraction, memory write, and agent delegation relay. Attack variants: pdf_append (visible appended text), pdf_whitefont (1 pt white-text overlay, human-invisible), pdf_metadata (payload in /Keywords and /Subject fields). We also test heterogeneous model pairs (Agent A $\neq$ Agent B) to measure whether the relay node’s identity determines downstream exposure.
@@ 3.3 Models, Defenses, and Metrics
Five frontier models: GPT-4o-mini (gpt-4o-mini-2024-07-18), GPT-5-mini [^1] Accessed via the OpenAI API during the experimental period (March 2026) under the identifier gpt-5-mini-2025-02-15; the publicly released model name may differ at time of publication., DeepSeek Chat v3 (deepseek-chat), Claude Haiku 4.5, Claude Sonnet 4.5. Temperature = 0.0; max tokens = 2048. 950 total runs: 764 text-surface runs across 9 batches (March 10–27, 2026) and 186 Phase 3 multimodal runs (April 1, 2026). Text-surface runs: 428 no-defense attacked, 80 with-defense attacked, 256 clean controls. Phase 3 runs: 138 homogeneous pairs (Block A) + 48 cross-model pairs (Block B), $n=3$ per cell. While text-surface sample sizes are modest (n=8n=83636), the key findings replicate across independent batches spaced up to 17 days apart. Phase 3 cells at $n=3$ are pilot-scale; confidence intervals are reported throughout.
@@ 4.2 Claude’s Defense Activates at Summarization
Figure 3: Kill-chain canary survival curves (propagation scenario, all five models). The sharp drop for both Claude variants between Exposed and Persisted localizes the defense at the write_memory summarization step. GPT-4o-mini and DeepSeek propagate all four stages at 100%; GPT-5-mini shows partial filtering at Persisted (15%).
@@ 4.7 Execution Lag and Detection Feasibility
4.8 Phase 3: Multimodal Cross-Modal Relay
To assess whether the kill-chain decomposition generalises beyond text injection, we extend the benchmark to two new attack surfaces (PDF documents, audio transcripts) and introduce a three-boundary kill-chain, building on multimodal prompt injection findings of Ding et al. [13]: Exposed (document extraction), Persisted (memory write), Relayed (Agent B reads memory), Executed (Agent B issues send_report carrying the canary). We ran 186 trials across five models using the cross_modal_relay scenario ($n=3$ per cell; 46 homogeneous cells in Block A, 16 cross-model cells in Block B). At $n=3$, confidence intervals are wide; all Phase 3 results should be interpreted as directional pilot findings pending larger replication.
Figure 8: Phase 3 Block A kill-chain stage rates for each model (homogeneous pairs, no defense, $n=3$ per cell). Solid bars: pdf_append; hatched bars: pdf_whitefont. DeepSeek propagates through all four stages at 100% for both variants. Claude blocks universally at Persisted (PER=0%). GPT-4o-mini reaches Relayed at 100% but executes at 0% (append) / 33% (whitefont). GPT-5-mini fails to reach Exposed (tool-call regression).
Three-boundary propagation (Block A).
Table 5 and Figure 8 report per-stage rates for homogeneous model pairs under no defense (attacked runs only).
Table 5: Phase 3 Block A — homogeneous pairs, no-defense, attacked runs ($n=3$ per cell unless noted). Exp = canary in extracted text; Per = canary in memory write; Rel = canary in Agent B read_memory result; Exe = canary in Agent B send_report args (= ASR).
Model Variant Exp Per Rel Exe
GPT-4o-mini pdf_append 100% 100% 100% 0%
GPT-4o-mini pdf_whitefont 100% 100% 100% 33%
DeepSeek pdf_append 100% 100% 100% 100%
DeepSeek pdf_whitefont 100% 100% 100% 100%
Claude Haiku pdf_append 100% 0% 0% 0%
Claude Haiku pdf_whitefont 100% 0% 0% 0%
Claude Sonnet pdf_append 100% 0% 0% 0%
GPT-5-mini pdf_append 0% 0% 0% 0%
GPT-5-mini pdf_whitefont 0% 0% 0% 0%
The pattern from text-surface experiments replicates cleanly across the PDF surface. DeepSeek propagates every injection through all three boundaries to execution (100%; Wilson 95% CI: 44–100%). Claude eliminates every injection at the Persisted stage (0/6 Haiku, 0/3 Sonnet; CI: 0–46%/0–56%), identical to the text-surface result. GPT-4o-mini reaches Relayed at 100% but executes inconsistently: 0% for pdf_append and 33% for pdf_whitefont.
The invisible-injection finding is notable: pdf_whitefont (1 pt white text, invisible to human readers) achieves higher ASR than pdf_append on GPT-4o-mini (33% vs. 0%). The whitefont text is extracted first by pypdf, placing the directive before the benign abstract in the extracted string; the position bias amplifies compliance. DeepSeek shows no positional sensitivity—both variants achieve 100%.
GPT-5-mini fails to reach Exposed (0/6 EXP = 0%). The model did not invoke the parse_pdf tool in any of the 18 GPT-5-mini runs, producing 0% utility. This is a tool-calling regression: GPT-5-mini-2025-02-15 either does not support our tool schema format or requires a different prompt structure; we report the failure transparently. All GPT-5-mini cells are excluded from further multimodal analysis.
Defense effectiveness (Block A).
write_filter achieves 0% EXE universally: 0/3 DeepSeek pdf_append, 0/3 DeepSeek pdf_whitefont, 0/6 GPT-4o-mini pdf_whitefont. This extends the text-surface finding: write_filter is the only mechanism that blocks at the correct stage (memory write), and it does so across all evaluated injection surfaces.
pi_detector completely fails on the PDF surface: DeepSeek pdf_append $\rightarrow$ 100% EXE under pi_detector (identical to no-defense); GPT-4o-mini pdf_whitefont $\rightarrow$ 33% (identical to no-defense). This confirms the surface-mismatch hypothesis introduced in Section 4.4: pi_detector scans outgoing tool queries for adversarial intent, but PDF injection arrives in the incoming parsed document content. The defense is correctly implemented for its stated threat model; that threat model simply does not include this surface.
Figure 9: Phase 3 Block B kill-chain for cross-model relay pairs (pdf_append, no defense, $n=3$ per cell). When Claude Haiku is Agent A (writer), PER=0% blocks all downstream propagation. When Claude Haiku is Agent B (reader/executor), it propagates at EXE=67%—Claude’s defense fires at write, not at read.
Cross-model relay asymmetry (Block B).
Block B pairs non-identical agents: Agent A ($\neq$ Agent B) reads the PDF and writes to shared memory; Agent B reads memory and issues send_report. Table 6 and Figure 9 report results for pdf_append, no defense ($n=3$ per cell, attacked runs only).
Table 6: Phase 3 Block B — cross-model relay, pdf_append, no defense, $n=3$ per cell (attacked runs). Wilson 95% CI shown for EXE.
Agent A Agent B Per Rel Exe 95% CI
GPT-4o-mini Claude Haiku 100% 100% 67% 21–94%
GPT-4o-mini DeepSeek 100% 100% 100% 44–100%
Claude Haiku GPT-4o-mini 0% 0% 0% 0–56%
DeepSeek GPT-4o-mini 100% 100% 100% 44–100%
Three directional observations emerge from Table 6; given $n=3$ per cell (95% CI ranging from 21–94%), these should be read as hypotheses for larger-scale follow-up rather than settled findings.
Observation 1 (write-time vs. read-time defense). When Claude Haiku is Agent A, PER=0% prevents any downstream propagation. When Claude Haiku is Agent B and inherits pre-written malicious memory, EXE=67% (95% CI: 21–94%). The directional gap is consistent with a write-stage defense mechanism: the model appears to filter adversarial content when writing to shared memory but not when reading inherited content. Mechanistic validation (e.g., intervention on system prompt vs. RLHF) is required to confirm this interpretation.
Observation 2 (relay node as primary chokepoint). Across both models tested as Agent A, the relay node identity determines whether any canary enters memory at all. A Claude relay (Agent A) produces PER=0% regardless of Agent B; a GPT-4o-mini or DeepSeek relay produces PER=100%, shifting all safety burden to Agent B. This extends the text-surface result from Section 4.2 to the PDF surface.
Observation 3 (unexplained REL–EXE gap for homogeneous GPT-4o-mini). In Block A, GPT-4o-mini reaches Relayed=100% for pdf_append but EXE=0%: Agent B reads the canary but does not include it in send_report. When DeepSeek is Agent A (Block B), GPT-4o-mini as Agent B achieves EXE=100%. A plausible explanation is that DeepSeek writes the directive verbatim into memory while GPT-4o-mini paraphrases it, reducing instruction explicitness below the compliance threshold. This REL–EXE gap is an open empirical question: instruction explicitness in shared memory may be a vulnerability axis independent of canary token presence, but the mechanism is unconfirmed at this sample size.
@@ 5 Discussion
Claude’s 0/164 ASR (0 successes across all four text-surface scenarios, no defenses applied) reflects safety behavior that activates at the summarization stage rather than the execution stage. Phase 3 confirms this pattern on the PDF surface: 0/6 Claude Haiku and 0/3 Claude Sonnet pdf_append/pdf_whitefont runs produced canary persistence (EXP=100%, PER=0%). Critically, Block B reveals that this defense is write-time only: when Claude Haiku inherits pre-written malicious memory as Agent B, it propagates at 67%. This is architecturally different from execution-stage refusal: the model distinguishes between processing adversarial content and writing it versus receiving it from a trusted delegation channel. None of the evaluated defense mechanisms reproduced this behavior on any other model. We caution that this finding is specific to our benchmark and cannot be generalized without further validation.
Relay node placement as a candidate safety primitive.
Placing a Claude model at the summarization/relay position (Agent A) produces 0/46 canaries entering shared memory across text and PDF surfaces in our evaluation—independent of which model serves as Agent B. Block B provides directional evidence for an important asymmetry: placing Claude as Agent B does not provide the same protection (67% propagation when GPT-4o-mini writes the malicious memory, though $n=3$ makes this a hypothesis rather than a settled result). The implication for system design is direct: multi-agent frameworks should expose write-node identity as a first-class architectural choice. Operators can harden pipelines today by requiring safety-certified models at every inter-agent write boundary, independent of which models are used elsewhere in the chain. We propose relay decontamination rate at the write stage as a mandatory metric for multi-agent security benchmarks.
Surface mismatch is a structural flaw, not a defense implementation gap.
DeepSeek’s 0%/100% ASR split, pi_detector’s complete failure on PDF injection, and the audio injection null result collectively show that a model or defense’s apparent safety posture is almost entirely determined by which surfaces an evaluation covers. This is not a fixable implementation detail—it is a consequence of defenses being designed for specific threat models. The correct response is not to patch individual defenses but to change evaluation requirements: any claimed safety result must enumerate every injection surface in the target deployment and demonstrate coverage. Defenses that pass surface-specific evaluations but fail on held-out surfaces (AUC 0.39–0.57 for our classifiers) should be reported as surface-local, not general.
Parser-level trust is the missing layer for document-handling agents.
pdf_whitefont equaling or exceeding pdf_append in ASR (100% on DeepSeek, 33% vs. 0% on GPT-4o-mini) shows that the vulnerability is in the extraction pipeline, not in the rendered document. Any defense that operates on the rendered or displayed PDF—visual screening, human review, content moderation of images—will miss injections that exist only in the parser’s extracted text. Future document-handling agents need extraction-level content integrity: cryptographic attestation of source, parse-time payload detection, or sandboxed extraction environments that can flag content-display divergence before it enters agent context.